How many grams of diethyl ether are formed from .55 mol of ethanol


1
Expert's answer
2020-11-09T14:03:44-0500

At first, balanced chemical equation should be written:

2 C2H5OH = C2H5OC2H5 + H2O.

n(C2H5OC2H5)=n(C2H5OH)/2=0.55/2=0.275(mol)

m=n×M

M=4×12+10×1+16=74(g/mol)

m=0.275×74=20.35(g)